OpenStudy (some.random.cool.kid):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for two different years: Number of years 1 2 3 Option 1 (amount in dollars) 1300 1690 2197 Option 2 (amount in dollars) 1300 1600 1900 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. (2 points) Part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points)

OpenStudy (some.random.cool.kid):

@perl

OpenStudy (some.random.cool.kid):

@perl you there

OpenStudy (some.random.cool.kid):

ill be right back

OpenStudy (perl):

ok

OpenStudy (some.random.cool.kid):

alright im back

OpenStudy (some.random.cool.kid):

@PERL

OpenStudy (perl):

Number of years 1 2 3 Option 1 (amount in dollars) 1300, 1690, 2197 Option 2 (amount in dollars) 1300, 1600, 1900

OpenStudy (perl):

option 1 is exponential because it changes by a factor of 1.3 each year option 2 is linear because it increases by a fixed amount each year, here 300 .

OpenStudy (some.random.cool.kid):

hmmm

OpenStudy (some.random.cool.kid):

does that describe this ... or is there another step? "describe the value of the investment after a fixed number of years"

OpenStudy (some.random.cool.kid):

oh nvm i get it

OpenStudy (some.random.cool.kid):

but what about part B @perl

OpenStudy (some.random.cool.kid):

if that was part b then part a i need

OpenStudy (perl):

part b) is to find the linear and exponential functions

OpenStudy (some.random.cool.kid):

how would i solve that

OpenStudy (some.random.cool.kid):

@perl

OpenStudy (some.random.cool.kid):

@perl how do i solve part b for this

OpenStudy (some.random.cool.kid):

@ganeshie8

OpenStudy (some.random.cool.kid):

@wio

OpenStudy (some.random.cool.kid):

ok

OpenStudy (perl):

for part b) Option 1: Directions: Fit data to exponential model f(n) = a * b^n if you divide 1300 by 1.3 , you will get 1000. Therefore n=0 , f(0) = 1000 so the exponential function is f(n) = 1000 * (1.3)^n Option 2: Directions: Fit data to a linear model f(n) = m*n + b if you calculate slope you get m = 300 and subtract 300 to get n=0, f(0) = 1000 So the linear model is f(n) = 300*n + 1000

OpenStudy (some.random.cool.kid):

ok, i know its against the rules but since its 12:05 am for me im kind of going to need you to explain while giving the answer because i have to go to bed soon.

OpenStudy (some.random.cool.kid):

@perl

OpenStudy (some.random.cool.kid):

@ganeshie8

OpenStudy (some.random.cool.kid):

@ganeshie8

OpenStudy (some.random.cool.kid):

@wio

OpenStudy (some.random.cool.kid):

@DanJS

OpenStudy (some.random.cool.kid):

@DanJS do you know this stuff

OpenStudy (danjs):

yeah,

OpenStudy (danjs):

pearl did it

OpenStudy (some.random.cool.kid):

can you give me the answer i have to go plus ill repost this because its not posted correctly

OpenStudy (some.random.cool.kid):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for two different years: Number of years 1 2 3 Option 1 (amount in dollars) 1300 1690 2197 Option 2 (amount in dollars) 1300 1600 1900 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. (2 points) Part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points) Part C: Belinda wants to invest in an option that would help to increase her investment value by the greatest amount in 20 years. Will there be any significant difference in the value of Belinda's investment after 20 years if she uses option 2 over option 1? Explain your answer, and show the investment value after 20 years for each option. (4 points)

OpenStudy (some.random.cool.kid):

there

OpenStudy (danjs):

option 1: y(x) = A*b^x use 2 points to find, A and b

OpenStudy (danjs):

(0,1000) gives y(0) = 1000 = A*b^0 1000 = A The function is now y(x) = 1000 *b^x

OpenStudy (danjs):

(1 , 1300) second point y(1) = 1300 = 1000*b^1 1300 = 1000b b=1.3 The function is y(x) = 1000*(1.3)^x

OpenStudy (danjs):

The second is a line

OpenStudy (some.random.cool.kid):

ok which part is that

OpenStudy (danjs):

I just did option 1,

OpenStudy (danjs):

option 2 is a line, starts at y intercept of 1000 investment, and slope of 300 dollars per year y(x) = 300x + 1000

OpenStudy (danjs):

pearl already did all this , i just scrolled up there

OpenStudy (danjs):

Graph those 2 functions, and describe what you see

OpenStudy (some.random.cool.kid):

first question? what part did you do theres 3 parts second question perl start it but can you give me all the answers because i have to go its 12:30 am?

OpenStudy (danjs):

option1 y(x) = 1000*(1.3)^x option2 y(x) = 300x + 1000 Graph them, and describe what you see.

OpenStudy (some.random.cool.kid):

i dont have graph :/

OpenStudy (danjs):

part c, look at x=20, and go straight up, which one has a larger y value, it will be the exponential option 1.

OpenStudy (danjs):

desmos.com

OpenStudy (some.random.cool.kid):

dan815

OpenStudy (some.random.cool.kid):

@dan815

OpenStudy (some.random.cool.kid):

@Nnesha

OpenStudy (some.random.cool.kid):

ok im sorry but i really dont have time for this i need someone who can give me the answer now for part a part b and part c because its almost 1 A.M. and i have to go to bed...

OpenStudy (some.random.cool.kid):

@Abhisar

OpenStudy (some.random.cool.kid):

zepdrix

OpenStudy (some.random.cool.kid):

@zepdrix

OpenStudy (some.random.cool.kid):

@sammixboo

OpenStudy (some.random.cool.kid):

@sammixboo

OpenStudy (some.random.cool.kid):

@perl

OpenStudy (some.random.cool.kid):

@pooja195

OpenStudy (some.random.cool.kid):

i really dont have time for this i need someone who can give me the answer now for part a part b and part c because its almost 1 A.M. and i have to go to bed...

OpenStudy (some.random.cool.kid):

@blurbendy

OpenStudy (some.random.cool.kid):

@bibby

OpenStudy (some.random.cool.kid):

@ShadowLegendX

OpenStudy (some.random.cool.kid):

@sammixboo

OpenStudy (some.random.cool.kid):

@Nnesha

OpenStudy (some.random.cool.kid):

@wio

OpenStudy (some.random.cool.kid):

@dan815

OpenStudy (some.random.cool.kid):

@ganeshie8

OpenStudy (some.random.cool.kid):

@Hero

OpenStudy (some.random.cool.kid):

@nincompoop